OpenStudy (anonymous):

y varies inversely as three times x. When x = 3, y = 4. What is the value of x when y = 2? can someone help me solve?

hero (hero):

Diya! Diya!

OpenStudy (diyadiya):

y varies inversely as three times x \[y=k \times \frac{1}{3x}\] \[4=k \times \frac{1}{3*3}=\frac{k}{9}\]\[4= \frac{k}{9} \ \ or \ \ k= 9*4=36\]\[k=36\] \[ Now\ for \ y=2\] \[2= k \times \frac{1}{3x}= \frac{k}{3x}\]\[2= \frac{k}{3x}=\frac{36}{3x}\]\[2*3x=36 \]\[6x=36\]\[x=\frac{36}{6} =6\]\[x=6\]

hero (hero):

varies "inversely". Sometimes I post too quickly

OpenStudy (anonymous):

This is known as inverse proportion :)

hero (hero):

Diya deserves the medal, not me, lol

OpenStudy (diyadiya):

Thanks (: